# Seer: Facilitating Structured Reasoning and Explanation via Reinforcement Learning

### v1

@@ Appendix C Implementation Details
C.1 Stopping condition
For a fair comparison, we use T5-large model as the policy. However, we observe that due to the number of parameters as well as the issue about only a few "End" instances, T5-large tends to perform "Reason" actions more frequently. Moreover, unlike GPT-4, T5-large is less able to recognize when a hypothesis has been inferred and when to stop. Therefore, we attach other stopping conditions in addition to the "End" action: (1) The semantic similarity between the intermediate conclusion and the hypothesis exceeds a predefined threshold, i.e., BLEURT($i_{*}$, $h$) > 1. (2) Exceeding the maximum number of steps (set to 20 in this paper).
@@ C.2 Alignment algorithm
Following (Dalvi et al., 2021), we evaluate the intermediate steps based on Jaccard similarity. Specifically, the intermediate nodes $i_{*}$ in $T_{\text{pred}}$ are aligned with the intermediate nodes in $T_{\text{gold}}$ that have the maximum Jaccard similarity. If the Jaccard similarity between the intermediate node in $T_{\text{pred}}$ and all intermediate nodes in $T_{\text{gold}}$ is zero, it is aligned with "NULL". Note that only the intermediate node that is perfectly matched, i.e., the Jaccard similarity is 1, is considered as a correct step. Figure 5 provides a detailed illustration of this process. The alignment process is similar in the reasoning graphs (Ribeiro et al., 2023).
@@ C.3 Retriever for Task 3
In Task 3 of EntailmentBank, it is first necessary to retrieve relevant sentences from the corpus (Dalvi et al., 2021). The focus of this paper is to enhance the structured reasoning ability of the policy. Therefore, we adopt the retriever and its associated parameters from previous work (Hong et al., 2023), which is based on the all-mpnet-base-v2 model (Reimers and Gurevych, 2019). For a fair comparison, we retrieve the top 25 most relevant sentences as $X$ for Task 3.
@@ C.4 Entailment Module
The entailment module is also based on T5-large, taking premises as input and generating intermediate conclusions. Our primary focus is to enhance the structured reasoning ability of the policy through a reasonable RL method, such as, our proposed Seer in this paper. Therefore, similar to the retriever, we employ the entailment module that has already been trained in previous work (Hong et al., 2023), which also aids in a fair comparison.
@@ C.6 Entailment Tree Construction
To evaluate the correctness of each step, we have to reconstruct the trajectory into an entailment tree $T_{\text{pred}}$ for comparison with $T_{\text{gold}}$. Figure 5 illustrates this process. We consider the last intermediate conclusion as the hypothesis and then construct the predicted entailment tree based on the reasoning relationship of each step (each step is equivalent to a subtree). The reconstruction process is similar in the reasoning graphs (Ribeiro et al., 2023).
@@ C.7 Running time
In our experimental environment, the average training time per entailment tree for our Seer is 6.98 seconds, and the average inference time per entailment tree for Seer is 3.91 seconds. The inference time per entailment tree for RLET (Liu et al., 2022) is 9.34 seconds. The inference time per entailment tree for FAME (Hong et al., 2023) is 30.77 seconds, as reported in their papers. FAME leverages Monte-Carlo planning, necessitating the exploration of numerous nodes to enhance the reasoning capability of the policy, thus requiring considerable computational time. Our proposed Seer significantly surpasses FAME in terms of both efficiency and effectiveness, representing substantial progress.
@@ C.8 Experiment Environments
All experiments were conducted on Ubuntu 22.04 equipped with NVIDIA A100 GPUs. Our code mainly depends on python 3.10444https://www.python.org/ and PyTorch 2.0.1555https://pytorch.org/. The pre-trained language models are derived from HuggingFace Transformers [^6] https://huggingface.co/.
@@ Appendix D Metrics Details
For EntailmentBank, we follow (Dalvi et al., 2021) and evaluate the entailment tree $T_{\text{pred}}$ with three dimensions:
$\bullet$ Leaves: To evaluate the leaf nodes of $T_{\text{pred}}$, we compute F1 by comparing $X_{\text{pred}}$ to $X_{\text{gold}}$.
$\bullet$ Steps: To evaluate the structural correctness of each step, we compare all steps between $T_{\text{pred}}$ and $T_{\text{gold}}$ and compute F1. A predicted step is considered structurally correct if its premises (e.g., $x_{*}$, $i_{*}$) exactly match the gold premises.
$\bullet$ Intermediates: To evaluate the intermediate conclusions, we compare the aligned intermediate conclusions and compute F1. A predicted intermediate conclusion is deemed correct if the BLEURT score (Sellam et al., 2020) exceeds 0.28.
For each dimension, AllCorrect score is 1 if F1 is 1, otherwise 0. Given the above scores, we employ the Overall AllCorrect metric to comprehensively evaluate $T_{\text{pred}}$, which takes a value of 1 if and only if all leaves, steps, and intermediates are all correct. Note that this is an extremely strict metric, where any deviations in $T_{\text{pred}}$ will result in a score of 0.
For the STREET benchmark, we follow (Ribeiro et al., 2023) to evaluate in terms of both the answer to the question and the quality of the reasoning graphs.
@@
$\bullet$ Reasoning Graph Accuracy: This metric compares the predicted reasoning graph and the golden reasoning graph from the aspects of the graph structure and the content of intermediate conclusions. Please note that this is a stringent metric, with minor deviations from the golden reasoning graph resulting in the prediction incorrect.
@@ Appendix E Illustrations and Case Study of Seer
Given a hypothesis $h$ and initial facts $X$, we first obtain the trajectory through the resoning process, as shown in Figure 4. The state update follows the markov decision process, meaning the current state only depends on the previous state. Figure 4 is an erroneous reasoning example to better illustrate the following steps. Then, we convert the trajectory $\tau$ into an entailment tree $T_{\text{pred}}$ and align it with $T_{\text{gold}}$ to assign reward for each intermediate conclusion (Figure 5). Furthermore, Figure 6 elucidates the issue of equivalent trajectories, and previous work can not accurately describe the logical relationship between different states in entailment trees.
@@
Figure 4: An illustration of the reasoning process of Seer. We start from the initial state $s_{1}$ where existing entailment steps $P_{1}=\varnothing$ and candidate sentences $C_{1}=X$. In each step, we sample an action and update the state until the reasoning is done. For the "Reason" action, we sent the premises to the entailment module. The new conclusion is added to the $C$, the premises is removed from $C$ and the entailment step is added to the $P$. For the "End" action, we end the reasoning process and output the trajectory.
@@
Figure 5: An illustration of the reward and alignment process of Seer. Each reasoning step is actually a subtree (the same in the reasoning graph (Ribeiro et al., 2023)). (1) We construct $T_{\text{pred}}$ using the last intermediate conclusion ($i_{4}$ in this example) as the hypothesis. (2) We calculate the Jaccard similarity between the intermediate node ($i_{*}$) in $T_{\text{pred}}$ and each golden intermediate node in $T_{\text{gold}}$ ($\hat{i}_{1}$ and $h$ in this example), and align with the maximum Jaccard similarity. In this example, $i_{1}$ is aligned with $\hat{i}_{1}$ due to $\text{JS}(i_{1},\hat{i}_{1})=1$. $i_{2}$ is aligned with "NULL". $i_{4}$ is aligned with $\hat{i}_{1}$ due to $\text{JS}(i_{4},\hat{i}_{1})=0.5$ and $\text{JS}(i_{4},h)=0.4$. (3) We assign rewards based on the alignment results. Note that $i_{3}$ ($s_{3}$) is a redundant step. $r_{1}=1$, $r_{2}=-1$, $r_{3}=-0.5$, and $r_{4}=-1$. Evidently, the reward for each state originates from the tree structure rather than chained trajectory. Therefore, the return of each state should also follow the tree structure (or graph structure in reasoning graphs) rather than chained trajectory.
@@
Figure 6: An illustration of the equivalent trajectory and the definition of return. As the reasoning steps of $s_{1}$, $s_{2}$, and $s_{3}$ are mutually independent, the execution order among these steps can be arbitrary. Thus, $\tau$ and $\tau^{\prime}$ are equivalent trajectories because they can be converted to the same entailment tree (Dalvi et al., 2021). As shown in blue box, previous work (Liu et al., 2022) defines the return (a.k.a expected cumulative reward) in a chained trajectory and would assign different returns to $s_{1}$ and $s_{2}$ in these equivalent trajectories. In contrast, as shown in red box, our structure-based return is defined based on the tree or graph structure inherent in structured reasoning, which is the same source of rewards. Our structure-based return will consistently allocate stable returns to equivalent trajectories, thereby promoting training stability and convergence. Furthermore, maintaining consistency between the sources of rewards and returns can significantly enhance the effectiveness of the policy.
@@ Appendix F Prompts for GPT-4
Figures 7 and 8 show the CoT (Wei et al., 2022) and ReAct (Yao et al., 2023b) prompts for GPT-4, and figures 9 and 10 show the prompts of thought generator and state evaluator in ToT (Yao et al., 2023a), respectively. We provide a detailed introduction to the task definition and guide the model to respond in the required format. We randomly selected three examples for in context learning. For a fair comparison, we use the same examples for CoT and ReAct, attributing similar thought to them. ReAct divides the dialogue into two rounds, "Thought" and "Action", to query GPT-4. For ToT, following (Yao et al., 2023a), we generate candidate actions using a thought generator and subsequently select and execute the optimal action through a state evaluator. Due to its exceptional reasoning capabilities and self-evaluation strategy, ToT achieves superior results compared to CoT and ReAct, as demonstrated in Table 2. However, ToT requires higher costs in comparison to CoT and ReAct.
@@ Appendix G Error Analysis
We conduct comprehensive error analysis on Task2 and Task3 of EntailmentBank.
@@ G.1 Error Analysis of Task2
(1) Reasoning Step Error (62%). This is the main source of errors, and predominantly depends on whether the policy can select the correct premise. We observe that a small portion of the errors (accounting for 12.9% of this error type) use all the gold leaves, but have errors in the combination order. Compared to other reasoning step errors, the policy identified the correct premise. For example, the gold steps is "$x_{24}\;\&\;x_{5}\rightarrow i_{1}$; $i_{1}\;\&\;x_{23}\rightarrow h$" and the error steps predicted by Seer is "$x_{23}\;\&\;x_{5}\rightarrow i_{1}$; $i_{1}\;\&\;x_{24}\rightarrow i_{2}$".
(2) Early Termination Error (18%). We observe that the reasoning process may terminate prematurely and the existing entailment steps are all correct. On one hand, T5-large outputs ’End’ prematurely, unlike GPT-4 which can accurately judge when to stop. On the other hand, the entailment module might erroneously infer a hypothesis, leading to premature termination.
(3) Intermediate Conclusion Error (10%). This error type is different from the above error (where the entailment module prematurely infers a hypothesis). Intermediate conclusion error denotes errors triggered by incorrect entailment in the intermediate conclusions, despite having both correct leaves and steps. For a fair comparison, we used the entailment module that has already been trained in previous work Hong et al. (2023). These errors can be mitigated by training a better entailment module. Furthermore, these errors are correct in the reasoning part, which is also the our focus in this paper.
(4) Imperfect Evaluation (10%). We discover that some trees deemed as invalid are in fact valid, indicating that current automated metrics underestimate the validity of the trees. The most common reason is that there are multiple valid ways to construct an entailment tree. For example, consider the structure of a gold tree: "$x_{1}\;\&\;x_{2}\;\&\;x_{3}\rightarrow h$" may be predicted as: "$x_{1}\;\&\;x_{2}\rightarrow i_{1}$; $i_{1}\;\&\;x_{3}\rightarrow i_{2}$".
@@ G.2 Error Analysis of Task3
(1) Missing Gold Leaves Error (58%). Missing gold leaves error refers to the situation where the gold leaves are not included in the facts $X$ retrieved from the corpus. This situation will inevitably lead to an error in the predicted entailment tree, regardless of how powerful the policy is. The bottleneck of this error lies in the retrieval model. For a fair comparison, we directly use the retrieval model that has already been trained in previous work (Hong et al., 2023).
(2) Reasoning Step Error (42%). This error is similar to those in Task 2.
Here we will not discuss other errors included in Task 2, as we found that a major bottleneck of Task 3 is how to retrieve more gold leaves. The focus of this paper is to enhance structured reasoning through a sensible RL approach. In future work, we will continue to consider how to integrate the retrieval module into RL, further improving performance in Task 3.

### v3

@@ Appendix C Implementation Details
C.1 Stopping criteria
For a fair comparison, we use the T5-large model to represent the policy. However, we observe that T5-large tends to perform "Reason" actions more frequently, which is caused by the smaller number of model parameters and the issue of having only a few "End" instances. Moreover, unlike GPT-4, T5-large is less able to recognize when a hypothesis has been inferred and when to stop. Therefore, we attach two extra stopping criteria in addition to the "End" action: (1) The semantic similarity between the intermediate conclusion and the hypothesis exceeds a predefined threshold, i.e., BLEURT($i_{*}$, $h$) > 1. (2) Exceeding the maximum number of reasoning steps (set to 20 in this paper).
@@ C.2 Alignment algorithm
Following (Dalvi et al., 2021), we evaluate the intermediate steps based on Jaccard similarity. Specifically, the intermediate nodes $i_{*}$ in $T_{\text{pred}}$ are aligned with the intermediate nodes in $T_{\text{gold}}$ that have the maximum Jaccard similarity. If the Jaccard similarity between the intermediate node in $T_{\text{pred}}$ and all intermediate nodes in $T_{\text{gold}}$ is zero, it is aligned with "NULL". Note that only the intermediate node that is perfectly matched with a node in $T_{\text{gold}}$, i.e., the Jaccard similarity is 1, is considered as a correct step. Figure 5 provides a detailed illustration of this process. The alignment process is similar in the reasoning graphs (Ribeiro et al., 2023).
@@ C.3 Retriever for Task 3
In Task 3 of EntailmentBank, first, it is necessary to retrieve relevant sentences from the corpus (Dalvi et al., 2021). The research focus of this paper is to enhance the structured reasoning ability of the policy. Therefore, we directly adopt the retriever and its associated parameters proposed in previous work (Hong et al., 2023), which is based on the all-mpnet-base-v2 model (Reimers and Gurevych, 2019). For a fair comparison, we retrieve the top 25 most relevant sentences as $X$ for Task 3.
@@ C.4 Entailment Module
The entailment module is also based on the T5-large model, taking premises as input and generating intermediate conclusions. Our primary focus is to enhance the structured reasoning ability of the policy through RL, therefore, we directly employ the entailment module that has already been trained in previous work (Hong et al., 2023), which also aids in a fair comparison.
@@ C.6 Entailment Tree Construction
To evaluate the correctness of each reasoning step, we have to reconstruct the trajectory into an entailment tree $T_{\text{pred}}$ and compare it with $T_{\text{gold}}$. Figure 5 illustrates this reconstruction process. We consider the last intermediate conclusion as the hypothesis and then construct the predicted entailment tree based on the reasoning relationship of each step. The reconstruction process is similar in the reasoning graphs (Ribeiro et al., 2023).
@@ C.7 Running time
In our experimental setting, the average training time per entailment tree in Seer is 6.98 seconds, and the average inference time per entailment tree in Seer is 3.91 seconds. As reported in their papers, the inference time per entailment tree in RLET (Liu et al., 2022) and FAME (Hong et al., 2023) are 9.34 seconds and 30.77 seconds, respectively. FAME leverages Monte-Carlo planning, necessitating the exploration of numerous nodes to enhance the reasoning capability of the policy, thus requiring considerable computational time. Our proposed Seer significantly surpasses FAME in terms of both efficiency and effectiveness.
@@ C.8 Experiment Environments
All experiments were conducted on Ubuntu 22.04 equipped with NVIDIA A100 GPUs. Our code mainly depends on python 3.10555https://www.python.org/ and PyTorch 2.0.1666https://pytorch.org/. The pre-trained language models are derived from HuggingFace Transformers [^7] https://huggingface.co/.
C.9 Details of Reasoning Graphs
For the reasoning graphs in the STREET Benchmark, the implementation details are slightly different from the entailment trees. In the reasoning graphs, reasoning steps may possess multiple parent nodes, and a fact ($x_{*}$) or intermediate conclusion ($i_{*}$) may be utilized multiple times (Ribeiro et al., 2023). Therefore, in the reasoning graph, we refrain from incorporating previously used premises into $U_{t}$, instead continually expanding the candidate sentence set $C_{t}$ through newly derived intermediate conclusions. In other words, the state in the reasoning graphs is updated according to the following rules: $P_{t+1}=P_{t}\cup\{\text{\textless premises\textgreater}\rightarrow i_{t}\}$, $C_{t+1}=\{X\cup I_{t+1}\}$, and $I_{t+1}=I_{t}\cup\{i_{t}\}$.
C.10 Other Implementation Details
For GPT-4, we set the temperature to 0.7. For Tree of Thought, we set $b=5$ candidates at each step, and then vote to select the optimal action. Details regarding the prompts of CoT, ToT, and ReAct can be found in Appendix F. For all baselines, we obtain the optimal results based on experimental results or hyperparameter settings derived from the original papers. For our method, we initialize the critic with the encoder of the warm-up policy to expedite the convergence of the critic and facilitate policy updates. The hidden layer dimension of the MLP in the critic is set to 512.
@@ Appendix D Metrics Details
For EntailmentBank, we follow (Dalvi et al., 2021) and evaluate the entailment tree $T_{\text{pred}}$ using three dimensions:
$\bullet$ Leaves: To evaluate the leaf nodes of $T_{\text{pred}}$, we compute F1 by comparing $X_{\text{pred}}$ with $X_{\text{gold}}$.
$\bullet$ Steps: To evaluate the structural correctness of each step, we compare all steps between $T_{\text{pred}}$ and $T_{\text{gold}}$ and then compute F1. A predicted step is considered structurally correct if its premises (e.g., $x_{*}$, $i_{*}$) exactly match the gold premises.
$\bullet$ Intermediates: To evaluate the intermediate conclusions, we compare the aligned intermediate conclusions and then compute F1. A predicted intermediate conclusion is deemed correct if the BLEURT score (Sellam et al., 2020) exceeds 0.28.
For each dimension, the AllCorrect score is 1 if F1 is 1, otherwise 0. Given the above scores, we employ the Overall AllCorrect metric to comprehensively evaluate $T_{\text{pred}}$, which takes a value of 1 if and only if all leaves, steps, and intermediates are correct. Note that this is an extremely strict metric, where any deviations in $T_{\text{pred}}$ will result in a score of 0.
For the STREET benchmark, we follow (Ribeiro et al., 2023) and adopt two metrics, namely, the answer to the question and the quality of the reasoning graphs, to evaluate different methods.
@@
$\bullet$ Reasoning Graph Accuracy: This metric compares the predicted reasoning graph and the golden reasoning graph from the aspects of the graph structure and the content of intermediate conclusions. Please note that this is a stringent metric, with minor deviations from the golden reasoning graph resulting in the prediction being incorrect.
@@ Appendix E Illustrations and Case Study of Seer
Given a hypothesis $h$ and initial facts $X$, we first obtain the trajectory through the reasoning process, as shown in Figure 4. The state update follows the Markov decision process (Bellman, 1957), meaning the current state only depends on the previous state. Figure 4 is an erroneous reasoning example to better illustrate the following steps. Then, we convert the trajectory $\tau$ into an entailment tree $T_{\text{pred}}$ and align it with $T_{\text{gold}}$ to assign reward for each intermediate conclusion (as presented in Figure 5). Furthermore, Figure 6 elucidates the issue of equivalent trajectories, and previous work can not accurately describe the logical relationship between different states in entailment trees.
@@
Figure 4: An illustration of the reasoning process of Seer. Note that $a_{1}$ is a correct step, $a_{2}$ and $a_{4}$ are erroneous steps, and $a_{3}$ is a redundant step. We start from the initial state $s_{1}$ where existing entailment steps $P_{1}=\varnothing$ and candidate sentences $C_{1}=X$. In each step, we sample an action and update the state until the reasoning is done. For the "Reason" action, we sent the premises to the entailment module. The new conclusion is added to the $C$, the premises is removed from $C$ and the entailment step is added to the $P$. For the "End" action, we end the reasoning process and output the trajectory.
@@
Figure 5: An illustration of the reward and alignment process of Seer. Each reasoning step is a subtree (similarly, each reasoning step is a subgraph in the reasoning graph (Ribeiro et al., 2023)). (1) We construct $T_{\text{pred}}$ using the last intermediate conclusion ($i_{4}$ in this example) as the hypothesis. (2) We calculate the Jaccard similarity between the intermediate node ($i_{*}$) in $T_{\text{pred}}$ and each golden intermediate node in $T_{\text{gold}}$ ($\hat{i}_{1}$ and $h$ in this example), and align with the maximum Jaccard similarity. In this example, $i_{1}$ is aligned with $\hat{i}_{1}$ due to $\text{JS}(i_{1},\hat{i}_{1})=1$. $i_{2}$ is aligned with "NULL". $i_{4}$ is aligned with $\hat{i}_{1}$ due to $\text{JS}(i_{4},\hat{i}_{1})=0.5$ and $\text{JS}(i_{4},h)=0.4$. (3) We assign rewards based on the alignment results. Note that $i_{3}$ ($s_{3}$) is a redundant step. $r_{1}=1$, $r_{2}=-1$, $r_{3}=-0.5$, and $r_{4}=-1$. The reward for each state originates from the tree structure rather than the chained trajectory. Therefore, the return of each state should also follow the tree structure (or graph structure in reasoning graphs) rather than the chained trajectory.
@@
Figure 6: An illustration of the equivalent trajectory and the definition of return. As the reasoning steps of $s_{1}$, $s_{2}$, and $s_{3}$ are mutually independent, the execution order among these steps can be arbitrary. Thus, $\tau$ and $\tau^{\prime}$ are equivalent trajectories because they can be converted to the same entailment tree (Dalvi et al., 2021). As shown in blue box, previous work (Liu et al., 2022) defines the return (a.k.a cumulative reward) in a chained trajectory and would assign different returns to $s_{1}$ and $s_{2}$ in these equivalent trajectories. In contrast, as shown in red box, our structure-based return is defined based on the tree or graph structure inherent in structured reasoning, which is the same source of rewards. Our structure-based return will consistently allocate stable returns to equivalent trajectories, thereby promoting training stability and convergence. Furthermore, maintaining consistency between the sources of rewards and returns can significantly enhance the effectiveness of the policy.
@@ Appendix F Prompts for GPT-4
Figures 7 and 8 show the Chain-of-Thought (CoT) (Wei et al., 2022) and ReAct (Yao et al., 2023b) prompts for GPT-4, and figures 9 and 10 show the prompts of thought generator and state evaluator in Tree of Thought (ToT) (Yao et al., 2023a), respectively. We provide a detailed introduction to the task definition and guide the model to respond in the required format. We randomly selected three examples for in-context learning. For a fair comparison, we use the same examples for CoT and ReAct, attributing similar thoughts to them. ReAct divides the dialogue into two rounds, "Thought" and "Action", to query GPT-4. For ToT, following (Yao et al., 2023a), we generate candidate actions using a thought generator and subsequently select and execute the optimal action through a state evaluator. Due to its exceptional reasoning capabilities and self-evaluation strategy, ToT achieves superior results compared to CoT and ReAct, as shown in Table 2. However, ToT requires higher costs in comparison to CoT and ReAct.
@@ Appendix G Error Analysis
We conduct a comprehensive error analysis on Task2 and Task3 of EntailmentBank.
@@ G.1 Error Analysis of Task2
(1) Reasoning Step Error (62%). This is the main source of errors and predominantly depends on whether the policy can select the correct premise. We observe that a small portion of the errors (accounting for 12.9% of this error type) use all the gold leaves, but have errors in the combination order. Compared to other reasoning step errors, the policy identified the correct premise. For example, the gold steps are "$x_{24}\;\&\;x_{5}\rightarrow i_{1}$; $i_{1}\;\&\;x_{23}\rightarrow h$" and the error steps predicted by Seer are "$x_{23}\;\&\;x_{5}\rightarrow i_{1}$; $i_{1}\;\&\;x_{24}\rightarrow i_{2}$".
(2) Early Termination Error (18%). We observe that the reasoning process may terminate prematurely and the existing entailment steps are all correct. On one hand, T5-large outputs “End” prematurely, unlike GPT-4 which can accurately judge when to stop. On the other hand, the entailment module might erroneously infer a hypothesis, leading to premature termination.
(3) Intermediate Conclusion Error (10%). This error type is different from the above error (where the entailment module prematurely infers a hypothesis). Intermediate conclusion error denotes errors triggered by incorrect entailment in the intermediate conclusions, despite having correct leaves and steps. For a fair comparison, we used the entailment module that has already been trained in previous work Hong et al. (2023). It is noted that the reasoning part, which is the focus of our paper, is completely correct in this type of error, and this type of error can be mitigated by training a better entailment module.
(4) Imperfect Evaluation (10%). We discover that some trees deemed as invalid are valid in fact, indicating that current automated metrics underestimate the validity of the trees. The most common reason is that there are multiple valid ways to construct an entailment tree. For example, consider the structure of a gold tree: "$x_{1}\;\&\;x_{2}\;\&\;x_{3}\rightarrow h$" may be predicted as: "$x_{1}\;\&\;x_{2}\rightarrow i_{1}$; $i_{1}\;\&\;x_{3}\rightarrow i_{2}$".
@@ G.2 Error Analysis of Task3
(1) Missing Gold Leaves Error (58%). Missing gold leaves error refers to the case where the gold leaves are not included in the facts $X$ retrieved from the corpus. This case will inevitably lead to an error in the predicted entailment tree, regardless of how powerful the policy is. The bottleneck of this error lies in the retrieval model. For a fair comparison, we directly use the retrieval model provided in previous work (Hong et al., 2023).
(2) Reasoning Errors (42%). The four error types described in G.1 account for 42% in Task3.
We also discovered that the reasoning graph contains similar error types as found in entailment trees, as they both belong to structured reasoning.